# GATE OpenING: A Comprehensive Benchmark for Judging Open-ended Interleaved Image-Text Generation

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have made significant strides in visual understanding and generation tasks. However, generating interleaved image-text content remains a challenge, which requires integrated multimodal understanding and generation abilities. While the progress in unified models offers new solutions, existing benchmarks are insufficient for evaluating these methods due to data size and diversity limitations. To bridge this gap, we introduce GATE OpenING (OpenING), a comprehensive benchmark comprising 5,400 high-quality human-annotated instances across 56 real-world tasks. OpenING covers diverse daily scenarios such as travel guide, design, and brainstorming, offering a robust platform for challenging interleaved generation methods. In addition, we present IntJudge, a judge model for evaluating open-ended multimodal generation methods. Trained with a novel data pipeline, our IntJudge achieves an agreement rate of 82. 42% with human judgments, outperforming GPT-based evaluators by 11.34%. Extensive experiments on OpenING reveal that current interleaved generation methods still have substantial room for improvement. Key findings on interleaved image-text generation are further presented to guide the development of next-generation models. The OpenING is open-sourced at https://opening.github.io.
@@ 1 Introduction
Building upon the remarkable understanding and generation capabilities of Large Language Models (LLMs) [1, 63, 66, 64], Multimodal LLMs (MLLMs) are making progress in various tasks [81, 42, 84, 5]. However, generating interleaved image-text content remains challenging [62, 70, 38], despite its important role in both research and applications (e.g., multimodal reasoning [11, 46], education [37, 17] and design [58, 34]). Since human brains can naturally combine visual and textual signals for more efficient information exchange [25, 31], achieving such integrated ability is crucial for advancing towards Artificial General Intelligence (AGI).
@@
Figure 2: OpenING benchmark consists of 23 meta-topics (inner ring) which are further categorized into 56 specific tasks (see the number of tasks on the outer ring and details in Appendix Sec. A.1). Examples showcase interleaved generation in eight representative domains.
@@
As shown in Fig.1, the emergence of unified models that combine understanding and generation abilities opens new possibilities for interleaved generation [77, 92]. However, the lack of reliable benchmarks for evaluating interleaved image-text generation remains an obstacle [70, 61]. Most existing benchmarks evaluate text or image outputs separately and do not capture the complexities of generating both simultaneously [44, 83, 82, 60]. Interleaved benchmarks like OpenLEAF [4] and InterleavedBench [43] are limited in size, scope, and query diversity. For example, InterleavedBench includes only 815 instances across 10 tasks sourced from public datasets such as VIST [32], ActivityNet [36], and WikiHow [80], which is not representative of real-world needs and faces the risk of data contamination.
@@
To fill the gap, we introduce OpenING, a comprehensive benchmark for evaluating open-ended interleaved generation. Unlike previous benchmarks, OpenING offers a broader set of real-world data and tasks (e.g., brainstorming, recommendations and content creation) derived from daily scenarios like fashion, cooking, and travel. As shown in Fig. 2 and Table 1, the curated OpenING includes 5,400 instances of multi-step interleaved image-text content across 23 meta-topics and 56 tasks, with diverse, carefully designed queries for various topics. To tackle the challenges of gathering and standardizing data from disparate domains, we develop an efficient annotation pipeline and produced high-quality human-annotated data, reducing data contamination risks.
@@
In addition, previous benchmarks often rely on GPT-based scoring metrics [4, 43], which are prone to be affected by the bias of GPT and potential data leakage in API uses [71]. To address the challenges of assessing open-ended multimodal generation, we introduce a robust judge model named IntJudge. We propose an Interleaved Arena to facilitate training data annotation and a Reference-Augmented Generation (RAG) approach to scale up the data size. Trained using the enhanced data pipeline, IntJudge achieves an 82.42% average agreement with human judgments, showing an 11.34% improvement over GPT-4o as a judge.
@@
We evaluate representative interleaved generation methods using our OpenING. Key findings from our experiments include: 1) Generating coherent and high-quality interleaved content remains challenging for all models, as human-annotated content receives the highest rating compared to the generated content; 2) While integrated pipelines (e.g. Gemini+Flux) outperforms end-to-end models (e.g. Anole) with more coherent image-text content and higher quality, end-to-end and two-stage generators (e.g. SEED-X) still hold great potential as the unified models continue to advance and can be combined with these methods; and 3) While text answers generated by GPT can be more informative than human-annotated answers, annotated natural images are still more favorable to generated images, highlighting the challenge of generating high-quality images. The major contributions of this paper are summarized as follows:
@@
  • A Comprehensive Leaderboard. We provide detailed rankings and analysis of interleaved generation methods and compare evaluations from our IntJudge and GPT-4o with human judgments. Our studies suggest that although current open-source end-to-end models perform less favorably than integrated generation pipelines, two-stage generators based on unified models deserve future exploration for the development of interleaved generation.
@@ 2 Related Work
Interleaved Image-Text Generation.
Development of MLLMs has greatly pushed forward interleaved image-text generation [35]. Early models like Stable Diffusion [53, 20], DALL-E [52], and autoregressive (AR) methods such as VAR [65] and Lumina-mGPT [41] focused on unidirectional tasks, including image understanding and text-to-image generation. Flamingo [2] was the first to introduce capabilities for interleaved image-text content. More recent models, such as MiniGPT-5 [88] and SEED-series [24, 23, 79], achieve interleaved generation by combining the AR-based text generation and diffusion-based visual generation. Native AR models such as Emu-3 [70] and Chameleon [62] offer a unified framework capable of generating and reasoning over mixed-modal documents. Anole [16] reproduces the image generation capability of Chameleon through efficient fine-tuning on interleaved image-text data. Meanwhile, benchmarks for the evaluation of interleaved image-text generation are still emerging. Previous works, such as OpenLEAF [4] and InterleavedBench [43] focused on a small set of scenarios and lacked the depth and breadth for real-world applications. To achieve a more reliable and holistic evaluation of interleaved generation, we propose OpenING – a comprehensive benchmark consisting of real-world scenarios.
Evaluation of Open-ended Multimodal Generation. Evaluating open-ended multimodal generation is inherently challenging, due to the requirement of assessing both visual and textual quality in open domain [4, 73]. Existing text generation metrics, such as BLEU [49] and ROUGE [39], lack the measure of visual quality or text-image coherence. Visual quality metrics like FID [30] and IS [54], on the other hand, lack the consideration of textual elements. Metrics like CLIPScore [29] can measure text-image alignment but are unable to fully evaluate the quality of open-ended interleaved content, where multiple correct answers exist. GPT-based scoring [86, 43] provides better measurements to assess the diversity and coherence of the interleaved outputs. However, GPT tends to be biased and favors the contents generated by its own [71, 6]. Human evaluation, though reliable, is not scalable due to its laborious nature. To bridge this gap, we introduce IntJudge, a judge model that is highly aligned with human judgments in evaluating the open-ended multimodal generation. To avoid the instability of subjective scores [89, 14], our IntJudge mainly evaluates models through pairwise comparisons in an arena manner.
@@ 3.1 Problem Definition
The task of interleaved image-text generation involves generating a sequence of text and images based on a given prompt. Each interleaved generation model (referred to as a multimodal agent) receives an input prompt $\mathbf{P}$, which can be text-only or include both texts and images. The multimodal agent outputs an interleaved image-text sequence: $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{N}]$, where $N$ is the number of steps. Each element $\mathbf{s}_{i}=<\mathbf{T}_{i},\mathbf{I}_{i}>$ in step $i$ consists of a text segment $\mathbf{T}_{i}$ and an image $\mathbf{I}_{i}$. Each $\mathbf{s}_{i}$ is generated based on the prompt $\mathbf{P}$ and all outputs history as $\mathbf{s}_{i}=f(\mathbf{P},\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{i-1})$, where $f$ denotes the the generation function of an agent. The objective is to find an optimal output sequence set $\mathbf{S}^{*}$:
@@
where $\mathbf{s}^{*}_{i}$ in each step is semantically consistent with the input prompt while the coherence throughout the entire sequence. The performance of an agent is evaluated based on how well the generated $\mathbf{S}$ meets predefined criteria.
@@ 3.2 Data Curation
Collecting and annotating interleaved image-text data is inherently challenging due to the scarcity of high-quality data. Moreover, it is difficult to gather multimodal data from disparate domains and pair them consistently. We created OpenING over three months, with nearly 50 people involved in an efficient pipeline, which is shown in Fig. 3.
@@
Data Collection and Annotation. We collected interleaved image-text data from more than 20 sources, including social media (e.g., REDnote [^1] https://www.xiaohongshu.com/explore) and YouTube [^2] https://www.youtube.com, search engine (e.g. Google [^3] https://www.google.com), and open-source platforms (e.g. OpenDataLab [28]), etc. (see Appendix Sec. A.2). We designed an efficient annotation process involving a team of 28 professional annotators, monitored by 14 data experts throughout the process to ensure the highest quality. The maximum number of steps per instance was limited to ten to ensure usability with context restraints. All queries and answers are annotated manually via our developed tool, IntLabel.
Data Filtering and Quality Control. We performed cross-checks with annotators and data experts to ensure high standards of consistency, relevance, and coherence for each instance. Each task was required to include diverse sources and topics. In a few cases where data acquisition posed challenges, annotators were instructed to supplement the dataset with content generated by GPT-4o [48] and Stable Diffusion XL [51]. To further enhance data quality, we implemented a set of exclusive protocols for filtering unqualified data, which is illustrated in Fig. 3. The qualified data are resupplied to each task to ensure a certain amount after filtering.
Data Processing. We conduct the post-processing to ensure the linguistic consistency of the benchmark. We utilized the GPT-4o API to translate the annotated Chinese text to English in several common tasks, followed by a double-check from data experts to verify the accuracy of translations. Additionally, for images containing Chinese characters, we implemented image translation [^4] https://github.com/zyddnys/manga-image-translator from Chinese to English. We finally adjusted prompts in each task for desired generation results, which are detailed in Appendix Sec. A.3.
@@ 4.1 Interleaved Arena
Evaluating open-ended interleaved image-text generation is challenging due to 1) the complexity of assessing multiple images and text, and 2) the open-ended nature of interleaved generation, where multiple valid answers can exist. Given that pairwise comparison is more stable than subjective scoring [14], we develop the Interleaved Arena, which supports pairwise evaluation using three pipelines: human judges, GPT-based judges, and the proposed IntJudge.
In the Interleaved Arena, interleaved outputs from agents on the OpenING Test Set are saved in a unified format. In each evaluation round, judges compare outputs from two anonymous agents and rate the interleaved outputs based on seven criteria: Correctness, Image-Text Coherency, Multi-step Consistency, Content Quality, Human Preference Alignment, Completeness, and Content Richness (see Appendix Sec. B.1 for more details). To balance evaluation reliability and efficiency, we propose a roulette matching algorithm to sample $E$ distinct battle pairs for each data instance:
Let $\mathcal{K}$ represent the set of tasks, where each task $k\in\mathcal{K}$ contains $D_{k}$ data instances. The set of agents in the arena is denoted as $\mathcal{M}$. For each task $k$, a permutation $\sigma_{k}\in A_{|\mathcal{M}|}$ is sampled by randomly shuffling the agent order, where $A_{|\mathcal{M}|}$ is the set of all permutations of agents in $\mathcal{M}$. The set of sampled battle pairs is given by:
@@
where $i=1,2,\ldots,\mathcal{D}_{k}$. Additional sampling rounds are conducted to obtain a total of $E$ distinct battle pairs for each data instance, where $E\leq|\mathcal{M}|(|\mathcal{M}|-1)/2$. To avoid duplication, we maintain a set $\mathcal{R}_{k,d}$ for the $d$-th round to track previously sampled pairs:
@@
For a newly sampled pair $\sigma_{k,d}(a)$ and $\sigma_{k,d}(b)$, we enforce:
@@
Under assumption of uniform distribution, we define the coverage time $T_{k}$ to ensure all agents are evaluated in task $k$:
@@
and the overall expected coverage time is given by:
@@ 4.2 Judge Pipelines
GPT-based Judge. To enable scalability, we employ GPT-4o to automate the evaluation process. The GPT-4o is prompted to analyze interleaved outputs and decide the winner for each battle pair. Moreover, we use an additional prompt to obtain the score breakdown and explanations. While this allows for a scalable evaluation with explainability, GPT-based judges still have a high error rate due to their prior bias and lack of alignment with human preferences. Besides, using GPT will bring concerns about privacy, data leakage, and cost.
IntJudge. To address issues in GPT-based evaluators, we propose IntJudge for more accurate evaluations and better alignment with human preferences. As an offline judge, IntJudge efficiently handles large-scale evaluations with consistent criteria, ensuring fair and reproducible results for benchmarking interleaved image-text generation. Specifically, we explored different MLLMs including InternLM-XComposer2.5 (InternLMX2.5) [85] and Qwen2-VL [68], ultimately choosing Qwen2-VL-7B as the foundational model for training IntJudge to achieve an optimal balance between efficiency and accuracy.

### v3

# OpenING: A Comprehensive Benchmark for Judging Open-ended Interleaved Image-Text Generation

@@ Abstract
Multimodal Large Language Models (MLLMs) have made significant strides in visual understanding and generation tasks. However, generating interleaved image-text content remains a challenge, which requires integrated multimodal understanding and generation abilities. While the progress in unified models offers new solutions, existing benchmarks are insufficient for evaluating these methods due to data size and diversity limitations. To bridge this gap, we introduce OpenING, a comprehensive benchmark comprising 5,400 high-quality human-annotated instances across 56 real-world tasks. OpenING covers diverse daily scenarios such as travel guide, design, and brainstorming, offering a robust platform for challenging interleaved generation methods. In addition, we present IntJudge, a judge model for evaluating open-ended multimodal generation methods. Trained with a novel data pipeline, our IntJudge achieves an agreement rate of 82.42% with human judgments, outperforming GPT-based evaluators by 11.34%. Extensive experiments on OpenING reveal that current interleaved generation methods still have substantial room for improvement. Key findings on interleaved image-text generation are further presented to guide the development of next-generation models.
@@ 1 Introduction
Building upon the remarkable understanding and generation capabilities of Large Language Models (LLMs) [1, 64, 67, 65], Multimodal LLMs (MLLMs) are making progress in various tasks [91, 84, 42, 87, 5]. However, generating interleaved image-text content remains challenging [63, 71, 37], despite its important role in both research and applications (e.g., multimodal reasoning [11, 46], education [36, 17] and design [59, 34]). Since human brains can naturally combine visual and textual signals for more efficient information exchange [25, 31], achieving such integrated ability is crucial for advancing towards Artificial General Intelligence (AGI).
@@
Figure 2: OpenING benchmark consists of 23 meta-topics (inner ring) which are further categorized into 56 specific tasks (see the number of tasks on the outer ring and details in Supplementary Materials). Examples showcase interleaved generation in eight representative domains.
@@
As shown in Fig. 1, the emergence of unified models that combine understanding and generation abilities opens up new possibilities for interleaved image-text generation [79, 96]. However, the lack of reliable benchmarks to evaluate interleaved generation remains an obstacle [71, 62]. Most existing benchmarks evaluate text or image output separately, failing to capture the intricacies of simultaneously generating both  [44, 86, 85, 61]. Interleaved benchmarks like OpenLEAF [4] and InterleavedBench [43] are limited in size, scope, and query diversity. For example, InterleavedBench includes only 815 instances across 10 tasks sourced from public datasets such as VIST [32] and WikiHow [83]. These benchmarks do not adequately reflect real-world needs and are vulnerable to data contamination [78].
@@
To fill the gap, we introduce OpenING, a comprehensive benchmark for evaluating open-ended interleaved generation. Unlike previous benchmarks, OpenING offers a broader set of real-world data and tasks (e.g., brainstorming, recommendations, and content creation) derived from daily scenarios like fashion, cooking, and travel. As shown in Fig. 2 and Table 1, the curated OpenING includes 5,400 instances of multi-step interleaved image-text content across 23 meta-topics and 56 tasks, with diverse, carefully designed queries for various topics. To address the challenges of collecting and standardizing data from disparate domains, we develop an efficient annotation pipeline and produce high-quality human-annotated data, reducing data contamination risks.
@@
In addition, many previous benchmarks rely on GPT-based scoring metrics [4, 43], which are prone to be affected by inherent biases in GPT models and potential data leakage in API usage [72]. To overcome the challenges of assessing open-ended multimodal generation, we introduce IntJudge, a robust judge model. We also propose an Interleaved Arena to facilitate annotation of training data and a Reference-Augmented Generation (RAG) approach to scale up the data size. Trained with this enhanced data pipeline, IntJudge achieves 82.42% average agreement with human judgments, an 11.34% improvement over GPT-4o as a judge.
@@
We evaluate representative interleaved generation methods using our OpenING. Key findings from our experiments include: 1) Generating coherent and high-quality interleaved content remains challenging for all models, whereas human-annotated content consistently receives the highest ratings over generated content; 2) Integrated pipelines (e.g., Gemini + Flux) outperform end-to-end models (e.g., Anole) in terms of image-text coherence and visual quality, possibly due to the more developed foundation models. End-to-end and two-stage generators (e.g., SEED-X) with unified architectures show great potential, especially as they continue to advance and potentially take advantage of developed methods; and 3) While text answers generated by GPT can be more informative than human-annotated answers, annotated natural images remain preferable to generated images, highlighting the challenges in high-quality image generation. The major contributions of this paper are summarized as follows:
@@
  • A Comprehensive Leaderboard. We provide detailed rankings and analysis of interleaved generation methods and compare IntJudge and GPT-4o evaluations with human judgments. Our findings indicate that while current open-source end-to-end models lag behind integrated pipelines, end-to-end and two-stage generators with unified architectures present great potential and warrant further explorations to advance interleaved image-text generation.
@@ 2 Related Work
Interleaved Image-Text Generation. The development of MLLMs has greatly advanced interleaved image-text generation [35]. Early models like Stable Diffusion [53, 20], DALL-E [52], and autoregressive (AR) methods (e.g., VAR [66] and Lumina-mGPT [41]) focused on unidirectional tasks, such as image understanding and text-to-image generation. Flamingo [2] was the first MLLM to process interleaved image-text content. More recent models, such as MiniGPT-5 [92] and SEED series [24, 23, 81], achieve interleaved generation by combining AR-based text generation and diffusion-based visual generation. Native AR models such as Emu-3 [71] and Chameleon [63] offer a unified framework to generate and reason over mixed-modal documents. Anole [16] reproduces the image generation capability of Chameleon through efficient fine-tuning on an interleaved image-text data. However, benchmarks for evaluating interleaved image-text generation remain in early stages. Previous works, such as OpenLEAF [4] and InterleavedBench [43] focused on a small set of topics and lack the depth and breadth for real-world applications. To achieve a more reliable and holistic evaluation of interleaved generation, we propose OpenING based on comprehensive real-world scenarios.
Evaluation of Open-ended Multimodal Generation. Evaluating open-ended multimodal generation is inherently challenging due to the need to assess both visual and textual quality in open domains [4, 74, 56]. Existing text generation metrics, such as BLEU [49] and ROUGE [39], fall short in measuring visual quality and text-image coherence. Conversely, visual quality metrics like FID [30] and IS [54] lack consideration of textual elements. Contrastive metrics, such as CLIPScore [29] can measure text-image alignment but fail to fully evaluate the quality of open-ended interleaved contents, where multiple correct answers may exist. GPT-based scoring [89, 43] provides improved measurements to assess the diversity and coherence of the interleaved outputs. However, GPT tends to be biased and favors the contents generated by itself [72, 6]. Human evaluation, although reliable, is not scalable due to its laborious nature. To close this gap, we introduce IntJudge, a judge model that is highly aligned with human judgments in evaluating the open-ended multimodal generation. To mitigate the instability of subjective scores [93, 14], our IntJudge evaluates models through pairwise comparisons in an arena-style framework [38].
@@ 3.1 Problem Definition
The task of interleaved image-text generation involves generating a sequence of text and images based on a given prompt. Each interleaved generation model (referred to as a multimodal agent) receives an input prompt $\mathbf{P}$, which can be text-only or include both text and images. The multimodal agent outputs an interleaved image-text sequence: $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{N}]$, where $N$ is the number of steps. Each element $\mathbf{s}_{i}=<\mathbf{T}_{i},\mathbf{I}_{i}>$ in step $i$ consists of a text segment $\mathbf{T}_{i}$ and an image $\mathbf{I}_{i}$. Each $\mathbf{s}_{i}$ is generated based on the prompt $\mathbf{P}$ and all outputs history as $\mathbf{s}_{i}=f(\mathbf{P},\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{i-1})$, where $f$ denotes the the generation function of an agent. The objective is to find an optimal output sequence set $\mathbf{S}^{*}$:
@@
where $\mathbf{s}^{*}_{i}$ in each step is semantically consistent with the input prompt while maintaining coherence throughout the entire sequence. The performance of an agent is evaluated based on how well the generated $\mathbf{S}$ meets predefined criteria.
@@ 3.2 Data Curation
Collecting and annotating interleaved image-text data is inherently challenging due to the scarcity of high-quality data. It is particularly difficult to gather and pair multimodal data from disparate domains and ensure consistency [82]. We created OpenING over three months, with nearly 50 people involved in an efficient pipeline, which is shown in Fig. 3(a).
@@
Data Collection and Annotation. Interleaved image-text data was collected from more than 20 sources, including social media (e.g., Rednote [^1] https://www.xiaohongshu.com), video sharing websites (e.g., YouTube [^2] https://www.youtube.com), search engines (e.g. Google [^3] https://www.google.com), and open dataset platforms (e.g. OpenDataLab [28]). The complete list of data sources is provided in the Supplementary Material. To ensure the highest data quality, a team of 28 professional annotators contributed under the supervision of 14 data experts. They performed efficient manual annotations using the IntLabel tool that we developed. Annotations were organized into a standard format, and each instance was limited to ten steps to avoid potential breaking of context constraints.
Data Filtering and Quality Control. We conducted cross-checks with annotators and data experts to ensure consistency, relevance, and coherence in each instance. Each task was required to include diverse sources and topics. In cases where data acquisition is complex, the annotators were instructed to supplement the data set with content generated by GPT-4o [48] and Stable Diffusion XL [51]. To further enhance data quality, exclusive protocols are proposed to filter unqualified data. The qualified data were then redistributed to each respective task to reach the required quantity.
Data Processing. Post-processings were carried out to ensure the linguistic consistency of our benchmark. GPT-4o API is used to translate the annotated Chinese text to English, followed by data experts reviewing the accuracy. We also implemented image translation [^4] https://github.com/zyddnys/manga-image-translator to convert any Chinese characters to English in images. Finally, the prompts were refined for each task to achieve desired generation results, as detailed in the Supplementary Materials.
@@ 4.1 Interleaved Arena
Evaluating open-ended interleaved image-text generation is challenging due to the complexity of assessing multiple images and text, as well as the open-ended nature of generation, where multiple valid answers exist. Given that pairwise comparison is more stable than subjective scoring [14], we introduce Interleaved Arena, on which pairwise assessments were conducted using three evaluators: human judges, GPT-based judges, and the proposed IntJudge.
In the Interleaved Arena, interleaved outputs from agents on the OpenING Test Set are saved in a unified format. In each evaluation round, judges compare outputs from two anonymous agents and rate the interleaved outputs based on seven criteria: Correctness, Image-Text Coherency, Multi-step Consistency, Content Quality, Human Preference Alignment, Completeness, and Content Richness (see supplementary materials for more details). To balance evaluation reliability and efficiency, we propose a roulette matching algorithm to sample $E$ distinct battle pairs for each data instance.
Let $\mathcal{K}$ represent the set of tasks, and $\mathcal{M}$ denote a set of Arena agents. Each task $k\in\mathcal{K}$ has $D_{k}$ data instances. A permutation $\sigma_{k}\in A_{|\mathcal{M}|}$ is sampled by randomly shuffling the agent order, where $A_{|\mathcal{M}|}$ is the set of all agent permutations. The set of sampled battle pairs is given by:
@@
where $i=1,2,\ldots,\mathcal{D}_{k}$. Additional sampling rounds may be performed to obtain $E$ distinct battle pairs for each data instance, where $E\leq|\mathcal{M}|(|\mathcal{M}|-1)/2$. To avoid duplications, a set $\mathcal{R}_{k,d}$ is maintained in the $d$-th round, which stores all unique pairs sampled in previous rounds:
@@
For the current pair $\sigma_{k,d}(a)$ and $\sigma_{k,d}(b)$, we enforce:
@@
Under the assumption of uniform distribution, we define the coverage time $T_{k}$ to ensure all agents are evaluated in task $k$:
@@
The overall expected coverage time is written as:
@@ 4.2 Judge Pipelines
GPT-based Judge. To enable scalability, we employ GPT-4o to automate the evaluation process. The GPT-4o is prompted to analyze interleaved outputs and decide the winner of each battle pair. Moreover, we use an additional prompt to obtain the score breakdown and explanations. While this strategy allows for scalable and explainable evaluation, GPT-based judges still have a high error rate due to their prior bias and lack of alignment with human preferences. GPT also raises privacy, data leakage, and cost concerns.
IntJudge. To address issues in GPT-based evaluators, we propose IntJudge to improve evaluation accuracy and alignment with human preferences. As an offline judge, IntJudge provides efficient large-scale evaluations with consistent criteria, ensuring fair and reproducible results for benchmarking interleaved image-text generation. Upon exploring several MLLMs including InternLM-XComposer2.5 (InternLMX2.5) [88] and Qwen2-VL [69], we select Qwen2-VL-7B as the foundation model for training IntJudge, achieving an optimal balance between efficiency and accuracy.